in case of practical of refraction what will be the angle of deviation if the angle of incidence is 10, 20, 30, 40, 50, 60, 70

1
Expert's answer
2020-08-18T10:08:13-0400

The angle of deviation is "\delta=\theta-\rho."

First, find the angle of refraction:


"\rho(\theta)=\text{arcsin}\frac{\text{sin}\theta}{n}."


With refraction index "n=1.6" we have the following data:

"\rho(10°)=6°,\space\delta(10°)=4°;\\
\rho(20°)=12°,\space\delta(20°)=8°;\\
\rho(30°)=18°,\space\delta(30°)=12°;\\
\rho(40°)=24°,\space\delta(40°)=16°;\\
\rho(50°)=29°,\space\delta(50°)=21°;\\
\rho(60°)=33°,\space\delta(60°)=27°;\\
\rho(70°)=36°,\space\delta(70°)=34°.\\"